undeadknight26 (undeadknight26):

If f(x) = log2 (x + 4), what is f^-1(3)? 0 2 4 8

undeadknight26 (undeadknight26):

@Luigi0210

undeadknight26 (undeadknight26):

@pooja195 @robtobey @agent0smith @surjithayer @sleepyjess @Somy @Gabylovesyou @ganeshie8 @Hero

OpenStudy (homework123456):

@undeadknight26 the it is 4 = 3 then you put it in a fraction and it would be 2

undeadknight26 (undeadknight26):

Alright thanks :)

undeadknight26 (undeadknight26):

Your sure its 2 and not 4 right O.o

OpenStudy (somy):

@Luigi0210 come hereee

OpenStudy (homework123456):

@undeadknight26 im 100% sure

OpenStudy (luigi0210):

If I remember correctly isn't \(\Large log_{2} (x+4) =c \) the same as \(\Large2^c=(x+4) \) ?

OpenStudy (luigi0210):

So the exponential form is the inverse. Meaning that since c=3 \(\Large 2^3=(x+4) \) Which is \(\Large 8=x+4 \), so \(\Large 4=x \) Plugging it into x would of given us: \(\Large 2^c=7 \) which I don't seee as an option so...

undeadknight26 (undeadknight26):

Its 4.../.- I shoulda went wif 4 -_-

OpenStudy (somy):

x.x im so dumb here

OpenStudy (homework123456):

i cant belive you took my medal away

undeadknight26 (undeadknight26):

LOL well your answer was wrong so ._.